What is the probability of rolling a 1 both times?

What is the probability of rolling a 1 both times?

since there is a 1/6 chance of rolling a 1, you need to multiply that every time you role again so 6 x 6 = 36. the odds of rolling a 1 twice are 1/36.

What is the probability of rolling one six-sided die and obtaining a 1 or 2?

What is the probability of rolling one six-sided die and obtaining the following numbers? Because 2 is only found on one side of a six-sided die, then there is a 1/6 chance of rolling a two.

What is the probability of rolling a 7 or 11 with two dice?

If you’re still in the ball game, your chance of getting the second roll you need for an 11 is again 16, so the total chance that you will roll an 11 is 13⋅16=118. Adding these two independent probabilities, the chance of rolling either a 7 or 11 is 16+118=29.

What is the probability of rolling a 6 twice?

16
The probability of rolling two sixes on two rolls is 16 as likely as one six in one roll. To make up for this, a pair of dice should be rolled six times for every one roll of a single die in order to get the same chance of a pair of sixes.

What is the probability of getting the same number on two dice?

1/6
The probability of both events occurring is 1 * 1/6.

How to calculate the probability of rolling a number on a die?

For the odds of rolling a specific number (6, for example) on a die, this gives: Probability = 1 ÷ 6 = 0.167. Probabilities are given as numbers between 0 (no chance) and 1 (certainty), but you can multiply this by 100 to get a percentage. So the chance of rolling a 6 on a single die is 16.7 percent.

What is the probability of getting a three on the fifth roll?

The probability of getting a three on the fifth roll is = 0.0804 X = the number of independent trials until the first success. You play a game of chance that you can either win or lose (there are no other possibilities) until you lose. Your probability of losing is p = 0.57. What is the probability that it takes five games until you lose?

How is the probability of getting a 6 calculated?

So to get a 6 when rolling a six-sided die, probability = 1 ÷ 6 = 0.167, or 16.7 percent chance. Independent probabilities are calculated using: So to get two 6s when rolling two dice, probability = 1/6 × 1/6 = 1/36 = 1 ÷ 36 = 0.0278, or 2.78 percent.

What is the probability of getting a 6 on a six sided die?

So to get a 6 when rolling a six-sided die, probability = 1 ÷ 6 = 0.167, or 16.7 percent chance. Independent probabilities are calculated using: Probability of both = Probability of outcome one × Probability of outcome two So to get two 6s when rolling two dice, probability = 1/6 × 1/6 = 1/36 = 1 ÷ 36 = 0.0278, or 2.78 percent.